# EM and XRM Connectomics Imaging and Experimental Metadata Standards

### v1

@@ 1 Introduction
Neuroimaging datasets are rapidly evolving in scale, which may unlock critical insights into neuronal circuit organization and neurodegenerative disease including connectopathies. Supported by large-scale investments, such as the from the National Institutes of Health (NIH) Brain Research Through Advancing Innovative Neurotechnologies (BRAIN) Initiative [1], this growth in scale of neuroimaging is not a niche development as a growing number of research teams and multi-institution teams are generating groundbreaking datasets.
@@
In particular, there have been major advances in EM and XRM/XNH connectomics dataset generation, which enable the creation of increasingly large connectomes (wiring diagrams) with individual neuron and synapse resolution. Recent efforts have included large scale volumes from multiple species and brain regions, with increasing spatial extent. This includes imaging of C. elegans [2, 3], Mus musculus cortex [4, 5, 6, 7, 8], Drosophila [9, 10, 11], and Homo sapiens [12]. Individual image volumes are reaching the scale of hundreds of terabytes or even petabytes of data. New datasets include co-registration to additional modalities [8] as well as imaging of multiple individuals at different developmental time points [3]. In addition to these emerging datasets, there is continued development in imaging techniques, processing algorithms, data storage solutions, and analysis techniques [13, 14, 11, 15, 16, 17]. This proliferation of datasets and techniques hold tremendous promise large-scale, cross-dataset discovery, but also presents considerable challenges.
@@
This growth in neuroimaging is driving the need for sophisticated data management and metadata creation strategies to support the production and dissemination of complex datasets [18, 19]. As a result, the field is seeing a significant increase in collaborative neuroinformatics efforts, that are aimed at effectively storing, processing, and sharing vast and intricate datasets. These platforms include, DANDI [20], BrainLife [21], BossDB [13], BrainCircuits.io [16], OpenNeuro [22], NEMAR [23], BIL [24], NEMO [25], and DABI [26]. These large-scale neuroinformatics efforts span a range of modalities, including neurophysiology, genomics, transcriptomics, and functional and structural neuroimaging. These datasets are most impactful when they can be accessed and integrated across datasets and even modalities. This has been recently demonstrated by the BRAIN Initiative Cell Consensus Network atlas [27] data ecosystem [28], which fused data from across archives to create a comprehensive understanding of cellular organization in mammalian cortex. This collaboration among scientific consortia demonstrates the real-world benefits of Findable, Accessible, Interoperable and Reusable (FAIR) principles for data management [18]. These emerging large-scale datasets are challenging to reproduce and reuse, thus a standardized framework is necessary to enable researchers to evaluate data across these datasets [29]. This also applies to emerging EM and XRM/XNH datasets, particularly as these datasets increase in size and complexity.
@@
Given the recent growth in EM and XRM/XNH connectomics datasets, there is now a presenting need for similar neuroinformatics efforts and standardization for connectomics data. To address the needs of the EM and XRM/XNH connectomics community, tools for data storage, visualization and analysis such as BossDB [13], DVID [17], webKnossos [30], BrainCircuits.io [16], cloud-volume [31], Neuroglancer [32], neuPrint [33], and CAVE [14] are being developed and adopted by EM and biology communities [34]. However, there has not yet been a comprehensive effort to standardize EM and XRM/XNH connectomics datasets and provide metadata guidelines for these tools.
@@
While data standards have been proposed for neurophysiology (Neurodata Without Borders, [20]), neuroimaging (BIDS, [35]), and 3D microscopy [36], challenges in data size, data types, and acquisition techniques limit the application of these standards to EM and XRM/XNH connectomics. To overcome these barriers, novel metadata standards must be developed while ensuring harmonization with existing standards for other modalities. To address this, the Big-Data Electron-Microscopy for Novel Community Hypotheses: Measuring And Retrieving Knowledge (BENCHMARK) team formed a working group to create image and experimental metadata standardization recommendations for the EM connectomics community. This document presents the BENCHMARK Image and Experimental Metadata Standards v1.1 for the connectomics community aiming to refine neuroimaging standards to support large-scale programs, archives, and tools. A companion paper describes standards for annotations and data products derived from these data.
@@ 2.2 Harmonization Efforts with Existing Standards and Datasets
  • Metadata services for the archives NEMO [25], DANDI [20], BossDB [13], DViD [17], webKnossos [30] and BIL [24]
  • The standard for 3D Microscopy [36]
  • Neurodata Without Borders Standard [20]
  • Brain Imaging Data Structure (BIDS) Standard [35]
  • Existing standards of the Volume EM community [37]
@@
Based on Working Group input and meetings, we have formulated v1.1 BENCHMARK standards for Image Data and Experimental Metadata. We encourage adoption of this framework by the community, and continued revision to support emerging efforts, needs, and new technologies. These standards are recommended primarily for data dissemination to create FAIR datasets [18] (for example on archives, consortia portals, or project websites), rather than prescribing a particular approach to data manipulation within an experimental pipeline. Working versions of the standard are maintained at https://github.com/aplbrain/benchmark-metadata.
@@ 3.1.1 General Desirable Properties of Image and Data Formats
  • Fundamental concepts of data sharding [38] should be followed
@@ 3.1.2 Image Formatting Options
  • OME-Zarr (version 2 of the Zarr specification) - to be updated to version 3 of the Zarr specification when appropriate [39]
  • Cloud-volume [31]
  • Neuroglancer precomputed [32]
  • N5 [40]
These formats are broadly supported in the community, and tools such as TensorStore [41] can be utilized to provide a uniform API for reading and writing multiple array formats, including OME-Zarr, N5, and Neuroglancer precomputed. If a tool/archive uses an internal format, one or more of the above should be supported for compliance with the BENCHMARK v1.1 standard. Key issues to address moving forward include tracking provenance of data files, with immutable logs, versioned releases, and tracking of changes. Additional considerations include label mapping from data files to annotation IDs, an ongoing area of coordination between the raw image formats and annotation metadata formats. Additionally, support for lossless and lossy compression needs to be standardized.
@@ 3.2 Experimental Metadata for Projects
Figure 2: Simplified data model for storage of data, derived from a project hosted at BossDB (https://bossdb.org/project/kuan_phelps2020). This uses an example dataset including samples from Drosophila in multiple coordinate frames [42] The full metadata structure can be seen at https://github.com/aplbrain/benchmark-metadata
Fields documented in imaging and metadata standard have 18 classes that are used to denote metadata for a given project. This includes the structure of data channels and experiments within a project (Fig. 2). The metadata fields allow values consisting of strings, integers, enumerations, floats or Booleans, and are essential elements to help define the projects fully. The classes provide fields that gives the field name, description of the field name, and allowed values. The metadata encompasses titles of projects, names of contributors, creators, URIs, and more (Table 1). This information can be vital to have when it comes to type of searches and queries, such as project types, specimens’ ages, sex, organizations involved with a particular project, and more. The full relationship of all classes, fields, and enumerations are shown in Fig. 3.
@@ 3.2.1 Project
The projects class contains 36 fields that give a high-level description of a dataset. These high-level descriptions include the title, short title, description, year, the DOI (Digital Object Identifier), and more (Table 1). The project entity is the highest-level object which gives an overview of the data available in a dataset. This enables users to reference a list of projects with a thorough description and query key project metadata. String identifiers allow for easy search, and embedded identifiers ensure data satisfies FAIR principles. Boolean values represent if a project is public or has a publication, and each project links to related papers and indicates any modification to the project. The project contains links to the collections which make up the project.
@@ 3.2.5 CoordinateFrame
The CoordinateFrame class consists of four entities specifying an extent for the X, Y, and Z axis of a three-dimensional volume. Each dimension is specified as a minimum and maximum float value in a list. The resolution of each voxel (in X, Y and Z) is specified for the CoordinateFrame. It is recommended, but not required, that CoordinateFrame objects share a common reference within a project. Future work will consider mappings to common coordinate frameworks, as they are established in the field.
@@ 3.3.3 LightMicroscopySpecific
This class is derived from the Standard metadata for 3D microscopy [36] and is included to promote compatibility with this standard. Future work will investigate further methods for co-registration and harmonization of these standards. The LightMicroscopySpecific class contains 28 fields to capture the properties of channels and experiments imaged with light microscopy techniques. This includes details on the microscope as well as experimental settings.
@@ 3.3.5 Taxonomy
The taxonomy class contains 8 fields that provide information to categorize and classify biological samples and species. These classes are Taxonomy ID, current name, GenBank common name, NCBI BLAST name, rank, genetic, mitochondrial genetic code, and common name. Taxonomy information can be found on the National Library of Medicine Taxonomy website [43].
@@ 3.3.6 ImageOrientation
This class is derived from the Standard metadata for 3D microscopy [36], and is used to capture information related to image orientation angles in LightMicroscopySpecific.
@@ 3.3.7 Landmark
This class is derived from the Standard metadata for 3D microscopy [36], and is used to capture information related to anatomical landmarks in LightMicroscopySpecific.
@@ 3.4.1 Contributor
The contributor class consists of 11 fields that provide information about personnel associated with a specific project. This class gives information on the individual or group members who provided skills, knowledge, and resources for a particular project. The standard specifies the role that the individuals played towards the project, and each project is assigned a point of contact. There is also a minimum of one creator for each project. Out of the 11 fields, 6 fields in the contributor class are associated with digital identifiers. These identifiers include ORCID (Open Researcher and Contributor Identifier), RRID (Research Resource Identifiers), ISNI (International Standard Name Identifier), ROR (Research Organization Registry), and a few others (Table 1).
@@ 3.4.2 License
Within the license class, there are 4 fields that give information about the licenses which govern usage of the data for this project. This specifies the terms and agreements that protect the rights of a project. In the license class (Table 1), the fields that are required include the URI to the rights, identifier, and corresponding DOI. The project may be associated with multiple licenses, which reference individual datasets within the given project. For example, all public projects on the archive BossDB are currently licensed under the Creative Commons Attribution 4.0 International License [44].
@@ 4 Discussion
The proposed standards build on and incorporate existing standards efforts in the community, including existing BRAIN Initiative Archives [20, 13, 24, 25], deployed solutions from the EM connectomics community [17, 31, 32, 33, 14, 16], and existing standards efforts [20, 36, 35]. The incorporation of existing file formats and standards in the BENCHMARK recommendations demonstrates a pragmatic approach, acknowledging the importance of building upon established practices. This strategy not only promotes interoperability among various EM Connectomics archives and tools but also facilitate the seamless integration of these standards into the existing research infrastructure. By prioritizing compatibility, the BENCHMARK standards aim to reduce barriers to adoption, ensuring a smooth transition for researchers while enhancing the overall accessibility and utility of EM connectomics data. This will be particularly critical to enabling multi-modal analysis [28] and secondary comparative connectomics analysis [45].

### v2

@@ 1 Introduction
Neuroimaging datasets are rapidly evolving in scale, which may unlock critical insights into neuronal circuit organization and neurodegenerative disease including connectopathies. Supported by large-scale investments, such as the from the National Institutes of Health (NIH) Brain Research Through Advancing Innovative Neurotechnologies (BRAIN) Initiative [jorgenson2015brain], this growth in scale of neuroimaging is not a niche development as a growing number of research teams and multi-institution teams are generating groundbreaking datasets.
@@
In particular, there have been major advances in EM and XRM/XNH connectomics dataset generation, which enable the creation of increasingly large connectomes (wiring diagrams) with individual neuron and synapse resolution. Recent efforts have included large scale volumes from multiple species and brain regions, with increasing spatial extent. This includes imaging of C. elegans [white1986structure, witvliet2021connectomes], Mus musculus cortex [kasthuri2015saturated, bock2011network, dorkenwald2022binary, schneider2020chandelier, microns2021functional], Drosophila [takemura2013visual, xu2020connectome, dorkenwald2023neuronal], and Homo sapiens [shapson2021connectomic]. Individual image volumes are reaching the scale of hundreds of terabytes or even petabytes of data. New datasets include co-registration to additional modalities [microns2021functional] as well as imaging of multiple individuals at different developmental time points [witvliet2021connectomes]. In addition to these emerging datasets, there is continued development in imaging techniques, processing algorithms, data storage solutions, and analysis techniques [hider2022brain, dorkenwald2023cave, dorkenwald2023neuronal, turner2022reconstruction, braincircuits, katz2019dvid]. This proliferation of datasets and techniques hold tremendous promise large-scale, cross-dataset discovery, but also presents considerable challenges.
@@
This growth in neuroimaging is driving the need for sophisticated data management and metadata creation strategies to support the production and dissemination of complex datasets [wilkinson2016fair, johnson2023maturity]. As a result, the field is seeing a significant increase in collaborative neuroinformatics efforts, that are aimed at effectively storing, processing, and sharing vast and intricate datasets. These platforms include, DANDI [rubel2022neurodata], BrainLife [hayashi2023brainlife], BossDB [hider2022brain], BrainCircuits.io [braincircuits], OpenNeuro [markiewicz2021openneuro], NEMAR [delorme2022nemar], BIL [benninger2020cyberinfrastructure], NEMO [ament2023neuroscience], and DABI [duncan2023data]. These large-scale neuroinformatics efforts span a range of modalities, including neurophysiology, genomics, transcriptomics, and functional and structural neuroimaging. These datasets are most impactful when they can be accessed and integrated across datasets and even modalities. This has been recently demonstrated by the BRAIN Initiative Cell Consensus Network atlas [network2021multimodal] data ecosystem [hawrylycz2023guide], which fused data from across archives to create a comprehensive understanding of cellular organization in mammalian cortex. This collaboration among scientific consortia demonstrates the real-world benefits of Findable, Accessible, Interoperable and Reusable (FAIR) principles for data management [wilkinson2016fair]. These emerging large-scale datasets are challenging to reproduce and reuse, thus a standardized framework is necessary to enable researchers to evaluate data across these datasets [megabrain]. This also applies to emerging EM and XRM/XNH datasets, particularly as these datasets increase in size and complexity.
@@
Given the recent growth in EM and XRM/XNH connectomics datasets, there is now a presenting need for similar neuroinformatics efforts and standardization for connectomics data. To address the needs of the EM and XRM/XNH connectomics community, tools for data storage, visualization and analysis such as BossDB [hider2022brain], DVID [katz2019dvid], webKnossos [boergens2017webknossos], BrainCircuits.io [braincircuits], cloud-volume [cloudvolume], Neuroglancer [neuroglancer], neuPrint [clements2020neu], and CAVE [dorkenwald2023cave] are being developed and adopted by EM and biology communities [collinson2023volume]. However, there has not yet been a comprehensive effort to standardize EM and XRM/XNH connectomics datasets and provide metadata guidelines for these tools.
@@
While data standards have been proposed for neurophysiology (Neurodata Without Borders, [rubel2022neurodata]), neuroimaging (BIDS, [gorgolewski2016brain]), and 3D microscopy [ropelewski2022standard], challenges in data size, data types, and acquisition techniques limit the application of these standards to EM and XRM/XNH connectomics. To overcome these barriers, novel metadata standards must be developed while ensuring harmonization with existing standards for other modalities. To address this, the Big-Data Electron-Microscopy for Novel Community Hypotheses: Measuring And Retrieving Knowledge (BENCHMARK) team formed a working group to create image and experimental metadata standardization recommendations for the EM connectomics community. This document presents the BENCHMARK Image and Experimental Metadata Standards v1.1 for the connectomics community aiming to refine neuroimaging standards to support large-scale programs, archives, and tools. A companion paper describes standards for annotations and data products derived from these data.
@@ 2.2 Harmonization Efforts with Existing Standards and Datasets
  • Metadata services for the archives NEMO [ament2023neuroscience], DANDI [rubel2022neurodata], BossDB [hider2022brain], DViD [katz2019dvid], webKnossos [boergens2017webknossos] and BIL [benninger2020cyberinfrastructure]
  • The standard for 3D Microscopy [ropelewski2022standard]
  • Neurodata Without Borders Standard [rubel2022neurodata]
  • Brain Imaging Data Structure (BIDS) Standard [gorgolewski2016brain]
  • Existing standards of the Volume EM community [iudin2023empiar]
@@
Based on Working Group input and meetings, we have formulated v1.1 BENCHMARK standards for Image Data and Experimental Metadata. We encourage adoption of this framework by the community, and continued revision to support emerging efforts, needs, and new technologies. These standards are recommended primarily for data dissemination to create FAIR datasets [wilkinson2016fair] (for example on archives, consortia portals, or project websites), rather than prescribing a particular approach to data manipulation within an experimental pipeline. Working versions of the standard are maintained at https://github.com/aplbrain/benchmark-metadata.
@@ 3.1.1 General Desirable Properties of Image and Data Formats
  • Fundamental concepts of data sharding [datasharding] should be followed
@@ 3.1.2 Image Formatting Options
  • OME-Zarr (version 2 of the Zarr specification) - to be updated to version 3 of the Zarr specification when appropriate [moore2023ome]
  • Cloud-volume [cloudvolume]
  • Neuroglancer precomputed [neuroglancer]
  • N5 [n5]
These formats are broadly supported in the community, and tools such as TensorStore [tensorstore] can be utilized to provide a uniform API for reading and writing multiple array formats, including OME-Zarr, N5, and Neuroglancer precomputed. If a tool/archive uses an internal format, one or more of the above should be supported for compliance with the BENCHMARK v1.1 standard. Key issues to address moving forward include tracking provenance of data files, with immutable logs, versioned releases, and tracking of changes. Additional considerations include label mapping from data files to annotation IDs, an ongoing area of coordination between the raw image formats and annotation metadata formats. Additionally, support for lossless and lossy compression needs to be standardized.
@@ 3.2 Experimental Metadata for Projects
Figure 2: Simplified data model for storage of data, derived from a project hosted at BossDB (https://bossdb.org/project/kuan_phelps2020). This uses an example dataset including samples from Drosophila in multiple coordinate frames [kuan2020dense] The full metadata structure can be seen at https://github.com/aplbrain/benchmark-metadata
Fields documented in imaging and metadata standard have 18 classes that are used to denote metadata for a given project. This includes the structure of data channels and experiments within a project (Fig. 2). The metadata fields allow values consisting of strings, integers, enumerations, floats or Booleans, and are essential elements to help define the projects fully. The classes provide fields that gives the field name, description of the field name, and allowed values. The metadata encompasses titles of projects, names of contributors, creators, URIs, and more (Table LABEL:tab:table1). This information can be vital to have when it comes to type of searches and queries, such as project types, specimens’ ages, sex, organizations involved with a particular project, and more. The full relationship of all classes, fields, and enumerations are shown in Fig. 3.
@@ 3.2.1 Project
The projects class contains 36 fields that give a high-level description of a dataset. These high-level descriptions include the title, short title, description, year, the DOI (Digital Object Identifier), and more (Table LABEL:tab:table1). The project entity is the highest-level object which gives an overview of the data available in a dataset. This enables users to reference a list of projects with a thorough description and query key project metadata. String identifiers allow for easy search, and embedded identifiers ensure data satisfies FAIR principles. Boolean values represent if a project is public or has a publication, and each project links to related papers and indicates any modification to the project. The project contains links to the collections which make up the project.
@@ 3.2.5 CoordinateFrame
The CoordinateFrame class consists of four entities specifying an extent for the X, Y, and Z axis of a three-dimensional volume. Each dimension is specified as a minimum and maximum float value in a list. The resolution of each voxel (in X, Y and Z) is specified for the CoordinateFrame in the VoxelSize field. This is an ImageResolution type which specifies the x,y, and z resolution along with the relevant unit. It is recommended, but not required, that CoordinateFrame objects share a common reference within a project. Future work will consider mappings to common coordinate frameworks, as they are established in the field.
@@ 3.3.3 LightMicroscopySpecific
This class is derived from the Standard metadata for 3D microscopy [ropelewski2022standard] and is included to promote compatibility with this standard. Future work will investigate further methods for co-registration and harmonization of these standards. The LightMicroscopySpecific class contains 28 fields to capture the properties of channels and experiments imaged with light microscopy techniques. This includes details on the microscope as well as experimental settings.
@@ 3.3.5 Taxonomy
The taxonomy class contains 8 fields that provide information to categorize and classify biological samples and species. These classes are Taxonomy ID, current name, GenBank common name, NCBI BLAST name, rank, genetic, mitochondrial genetic code, and common name. Taxonomy information can be found on the National Library of Medicine Taxonomy website [ncbitaxonomy].
@@ 3.3.6 ImageOrientation
This class is derived from the Standard metadata for 3D microscopy [ropelewski2022standard], and is used to capture information related to image orientation angles in LightMicroscopySpecific.
@@ 3.3.7 Landmark
This class is derived from the Standard metadata for 3D microscopy [ropelewski2022standard], and is used to capture information related to anatomical landmarks in LightMicroscopySpecific.
@@ 3.4.1 Contributor
The contributor class consists of 11 fields that provide information about personnel associated with a specific project. This class gives information on the individual or group members who provided skills, knowledge, and resources for a particular project. The standard specifies the role that the individuals played towards the project, and each project is assigned a point of contact. There is also a minimum of one creator for each project. Out of the 11 fields, 6 fields in the contributor class are associated with digital identifiers. These identifiers include ORCID (Open Researcher and Contributor Identifier), RRID (Research Resource Identifiers), ISNI (International Standard Name Identifier), ROR (Research Organization Registry), and a few others (Table LABEL:tab:table1).
@@ 3.4.2 License
Within the license class, there are 4 fields that give information about the licenses which govern usage of the data for this project. This specifies the terms and agreements that protect the rights of a project. In the license class (Table LABEL:tab:table1), the fields that are required include the URI to the rights, identifier, and corresponding DOI. The project may be associated with multiple licenses, which reference individual datasets within the given project. For example, all public projects on the archive BossDB are currently licensed under the Creative Commons Attribution 4.0 International License [creativecommons].
@@ 4 Discussion
The proposed standards build on and incorporate existing standards efforts in the community, including existing BRAIN Initiative Archives [rubel2022neurodata, hider2022brain, benninger2020cyberinfrastructure, ament2023neuroscience], deployed solutions from the EM connectomics community [katz2019dvid, cloudvolume, neuroglancer, clements2020neu, dorkenwald2023cave, braincircuits], and existing standards efforts [rubel2022neurodata, ropelewski2022standard, gorgolewski2016brain]. The incorporation of existing file formats and standards in the BENCHMARK recommendations demonstrates a pragmatic approach, acknowledging the importance of building upon established practices. This strategy not only promotes interoperability among various EM Connectomics archives and tools but also facilitate the seamless integration of these standards into the existing research infrastructure. By prioritizing compatibility, the BENCHMARK standards aim to reduce barriers to adoption, ensuring a smooth transition for researchers while enhancing the overall accessibility and utility of EM connectomics data. This will be particularly critical to enabling multi-modal analysis [hawrylycz2023guide] and secondary comparative connectomics analysis [barsotti2021neural].